# AI Assistants to Enhance and Exploit

### v1

@@ V-B Comparison with LLM baseline
Figure 6: Scoring comparison for LLMs.
@@ V-C Impact of reranking
Figure 7: Case Study 1: Comparison of RAG and reranking-enhanced RAG.
Figure 8: Case Study 2: Comparison of RAG and reranking-enhanced RAG.
@@
KSP can also be used to solve least squares
problems, using, for example, KSPLSQR...
@@
As described above, the option -info will print
information about the success of preallocation
during matrix assembly...
@@
V-D Inference latency
To evaluate inference latency, we measured separately the running time for the RAG process and the LLM response. This approach enabled us to quantify the overhead introduced by the RAG systems and assess their overall efficiency. As presented in Table II, adding the reranking stage increased the average RAG processing time by approximately $2.4$ times. However, the average running time for the reranking-enhanced RAG remained less than 11% of the average LLM response time.
@@ V-D Inference latency
TABLE II: Run time for RAG and the LLM (in seconds) on an Intel i7-11700KF 3.6GHz CPU.
VI Related Work
Recent advances in generative artificial intelligence, particularly in large language models, have significantly transformed software engineering. These models have demonstrated impressive capabilities in tasks such as code generation [18], bug detection [19], and automated code review [20]. Building on these capabilities, RAG systems [21] have extended the use of LLMs by integrating external knowledge sources during inference. Several studies ([22, 23]) have explored using RAG systems to support the development, maintenance, and transformation of legacy code bases.
@@ VI Related Work
VII Conclusion and Future Work
@@
Acknowledgment
We thank the PETSc developers for discussing this work as well as for their many contributions to the PETSc code base and community knowledge base, without which this work would not be possible.

### v2

@@ 5.2. Comparison with LLM baseline
Figure 6. Scoring comparison for LLMs.
@@ 5.3. Impact of reranking
Figure 7. Case Study 1: Comparison of RAG and reranking-enhanced RAG.
Figure 8. Case Study 2: Comparison of RAG and reranking-enhanced RAG.
@@
“KSP can also be used to solve least squares problems, using, for example, KSPLSQR …”
@@
“ As described above, the option -info will print information about the success of preallocation during matrix assembly …”
@@
5.4. Inference latency
To evaluate inference latency, we measured separately the running time for the RAG process and the LLM response. This approach enabled us to quantify the overhead introduced by the RAG systems and assess their overall efficiency. As presented in Table 2, adding the reranking stage increased the average RAG processing time by approximately $2.4$ times. However, the average running time for the reranking-enhanced RAG remained less than 11% of the average LLM response time.
@@ 5.4. Inference latency
Table 2. Run time for RAG and the LLM (in seconds) on an Intel i7-11700KF 3.6GHz CPU.
6. Related Work
Recent advances in generative artificial intelligence, particularly in large language models, have significantly transformed software engineering. These models have demonstrated impressive capabilities in tasks such as code generation (Chen et al., 2024; Nichols et al., 2024; Valero-Lara et al., 2023), HPC software development (Godoy et al., 2024), and compiler validation (Munley et al., 2024). Building on these capabilities, RAG systems (Lewis et al., 2020) have extended the use of LLMs by integrating external knowledge sources during inference. Several studies (Ali et al., 2024; Chen et al., 2023) have explored using RAG systems to support the development, maintenance, and transformation of legacy code bases.
@@ 6. Related Work
7. Conclusion and Future Work
@@
Acknowledgements.
We thank the PETSc developers for discussing this work as well as for their many contributions to the PETSc code base and community knowledge base, without which this work would not be possible. This material is based upon work supported by Laboratory Directed Research and Development (LDRD) funding from Argonne National Laboratory, provided by the Director, Office of Science, of the U.S. Department of Energy (DOE) under Contract No. DE-AC02-06CH11357. Research was partially supported by the U.S. DOE Office of Science Distinguished Scientist Fellows Program. This research used resources of the Argonne Leadership Computing Facility, a U.S. Department of Energy (DOE) Office of Science user facility at Argonne National Laboratory, and is based on research supported by the U.S. DOE Office of Science-Advanced Scientific Computing Research Program, under Contract No. DE-AC02-06CH11357.